OpenStudy (anonymous):

how to evaluate submission n=1 to infinity 2n+1/n^2 (n+1)^2

ganeshie8 (ganeshie8):

partial fractions + telescoping

hartnn (hartnn):

adjust the numerator 2n+1 = \( (n+1)^2 -n^2\)

hartnn (hartnn):

then separate out the denominator

OpenStudy (anonymous):

@ganeshie8 i have did the partial but i can't get the value for A and B

ganeshie8 (ganeshie8):

use hartnn's hint

ganeshie8 (ganeshie8):

\[\dfrac{2n+1}{n^2(n+1)^2} = \dfrac{(n+1)^2 - n^2}{n^2(n+1^2)} = \dfrac{1}{n^2} - \dfrac{1}{(n+1)^2} \]

ganeshie8 (ganeshie8):

telescope it now